OpenStudy (carlj0nes):

Prandtl Number does anybody know the relationship between α(thermal diffusivity) and k(thermal conductivity of a flowing fluid) Pr= (Molecular diffusivity of momentum)/(Molecular diffusivity of heat)=v/α=〖μc〗_p/k

OpenStudy (anonymous):

\[Pr = \frac{C_p \mu}{k}\] but \[\alpha = \frac{k }{\rho C_p}\]